# Generating coupled cluster code for modern distributed memory tensor software

### v1

@@ II THEORY AND IMPLEMENTATION
II.0.1 Coupled cluster amplitude equations and their generation
@@
where is the -th excitation, and the corresponding cluster amplitude. in our case denotes the Hartree Fock (HF) reference.
@@
We use and to denote occupied and virtual orbitals respectively, whereas denotes any orbitals.
After defining the similarity transformed normal-ordered Hamiltonian as , we can write the CC energy equations
@@
The explicit equation terms can then be derived using different techniques. We chose the diagrammatic scheme of Kállay and SurjánKállay and Surján (2001) (see the reasoning in section II.1). This general-order coupled cluster derivation scheme is based on a string representation of CC diagrams.
As depicted in Scheme 2, integer strings of length 13 is used to represent a diagram. Two diagrams are equivalent up to a sign if their strings are the same. There is a relatively simple algorithm that uses this representation to generate distinct diagrams for each excitation level of the CC Eqs. (3). Please refer to the original Ref. Kállay and Surján, 2001 for the full algorithm.
To find the explicit equation terms, tenpi translates the integer strings into line graph representation, i.e. an edge list of particle and hole lines connecting operators and external bra indices of equation Eq. (3). Such a representation makes the application of coupled cluster interpretation rules similar to when done by hand.
II.1 Design decisions
The initial goal was to create a programming environment for relativistic coupled cluster which separates science from the computational platform by getting tensor developments under the hood. This has three aspects:
@@ II.1 Design decisions
Automatic treatment of tensor symmetries, such as index-permutation symmetry and block-sparsity (for spatially-symmetric systems).
@@
The development was driven by practical considerations. The toolchain should not be restricted to a given flavor of the coupled cluster method, but should be written in a general way and be able to go to high orders of the theory. It should allow manipulations of the method by PhD and master students of chemistry. This led to the decision to base the formula generator on Kucharski-Bartlett diagrams,Shavitt and Bartlett (2009); Kucharski and Bartlett (1992) since diagrams are visual and thus more accessible for students than the other CC derivation schemes (see section I.1). The ease of use is one of the reasons why we chose python to implement the generator, aside from the fact that it is suitable for text processing and symbolic operations. The string representation of CC diagrams of Kállay and SurjánKállay and Surján (2001) was chosen because it can generate distinct connected diagrams and detect equivalent diagrams (see section II.0.1), both in an efficient and intuitive way.Kállay and Surján (2001)
@@
The diagram strings are next translated into directed graphs, by storing a list of edges (adjacency matrix). This representation makes the implementation of the CC interpretation rules readable. Advanced users can define a custom set of rules (customizability) or even extend the Diagram class to allow for multiple interaction operators and other generalizations, without having to change the rest of the code (modularity).
@@
To follow closely the hand-typeset style of the pedagogical texts of Refs. Crawford and Schaefer, 2000 and Shavitt and Bartlett, 2009, the toolkit prints the diagrams in PDF format using the CCDiag LaTeXpackage.Kats (2024) During the development of higher-order CC methods this visual representation proved itself invaluable for debugging.
@@
The flow of the tenpi code generator is depicted in Scheme 3. A key step is to optimize the expressions obtained. Since writing a global optimizer from scratch is beyond the scope of this work, we interfaced our code with OpMin, which was chosen due to its favorable performance for higher-order CC.Lai et al. (2012) However, for most tests, the FLOP-count based approach of OpMin generates too many intermediates. Higher-order CC tensors in our calculations are too large to be stored on disk due to prohibitive I/O costs. Our approach is to distribute them among the CPU and GPU RAM memory of the nodes. This comes with a limitation in the amount of available distributed memory due to communication overhead growing quadratically with the number of nodes for our simple round-robin distribution scheme. Therefore, we implemented our own secondary optimization, which rolls back part of the intermediates by plugging them back in their respective terms. The number of intermediates is reduced by about a factor of three at the expense of increasing the FLOP count (see Fig. 4). This approach is justified as long as the calculation stays communication-dominated. We do not reach a perfect balance, which would require setting up a communication-based cost-function. This is ongoing work.
@@
During the optimization, we originally represented contractions using the SymPy packageMeurer et al. (2017) but later we turned to the use of a custom representation of tensor contractions for greater flexibility. Even though SymPy is widely used, its API felt cumbersome for handling higher-order tensors.
@@
As we aim for scalability on modern GPU exascale machines, we turned to the use of a tensor library to execute the final tensor kernels.
II.2 Implementation
An equation term is implemented using the ‘Contraction’ class, which includes a list of tensors, a scalar factor and optionally permutation operators or a result tensor. Contraction class instances are collected in a list, which forms the left-hand side of Eqs. (3).
Tensor class includes a list of upper and lower indices. Each index is an Index class instance. Index class has its type (e.g. occupied, virtual) a convention according to which it is printed (e.g. ; ) and then a number correspoding to order in which the specific character comes in the class (e.g. 1, 2, 1). Types and conventions are easily customizable, which is important for the implementation of methods and symmetries with multiple index spaces or tiling.Mutlu et al. (2023) Equations can switch between conventions with a single command. This is important because tenpi supports multiple input and output formats to interface with ExaTENSOR, TAL-SH, OpMin, NumPy, including a possibility to input equations by hand. For the sake of modularity, all these libraries are integrated by extending one of two predefined interfaces: GetCodeInterface or ParseCodeInterface. These specify how to implement a new format of output or input respectively. Thanks to this design it is straightforward to interface tenpi with a new program.
@@ II.2 Implementation
| intermediates | 0.03 s | 40 s | 22 min | 7 h 10 min | 4 days 1 h |
@@
The derivation step of tenpi is very fast, see timings in Tab. 1. We have verified by comparing with Ref. Kállay and Surján, 2001 that tenpi generates the correct number of diagrams of the amplitude equation up to order 20. Even though it generates a highly-parallel code, tenpi itself is written as sequential python. Despite this, the most demanding step, global intermediate optimization, still performs quite well.
Aside from amplitude equations, tenpi can already generate matrix elements (c.f. SMITHShiozaki et al. (2008); Shiozaki (2017)) and is currently being extended to generate response density matrices.
During the development, the debugging of CCSDTQ proved quite difficult, due to a mistake in an algorithm generating permutations. A different, simple mistake appeared already at the CCSD level: For the CC diagram interpretation rules to yield a correct sign, e.g. for the case of amplitude-equation projected onto , the line starting in has to end in , and not in . This rule is often assumed implicitly.Shavitt and Bartlett (2009); Crawford and Schaefer (2000)
@@ III.1 Machines used
The calculations were performed on machines listed in Table 2 which includes their technical specification. When ExaTENSOR,Pototschnig et al. (2021) is applied for coupled cluster, the most relevant parameters for the performance are the interconnect bandwidth, RAM (random-access) and HBM (high-bandwidth) memory, and the specification of the GPU. Over the two supercomputer generations represented in the table the bandwidth and HBM have grown about fourfold, while RAM has basically stagnated. Despite the rapid development of GPU processing power, this shows that the evolution of other important parameters has been relatively slow over the last 5 years.
@@ III.2 Relativistic Hamiltonian
The state-of-the-art post-HF four-component relativistic calculations employ the no-pair approximation.Dyall and Faegri (2007) This approximation is well-substantiated for applications of chemical scale. Formally, it yields a Hamiltonian analogous to the non-relativistic case
where the indices run over the positive-energy spinors that span the one-electron basis.
In the present work we employ, unless otherwise stated, the exact two-component (X2C) Hamiltonian, XXX where the summations in Eq. (4) are limited to positive energy orbitals by construction.
Relativistic picture change corrections were added either using the AMFI package CCC or using the more recent amfX2C correctionKnecht et al. (2022).
@@
For the bechmarks, we used the UF6 molecule in dyall.v2z basisDyall (2023) with relativistic X2C Hamiltonian, and bond distance of Å. In the corresponding ground-state energy CC calculations, we used a (Dirac-)Hartree-Fock (HF) reference. Despite the high Oh symmetry, there is no related performance gain in the CC calculations as the implementation does not support it.
Furthermore, we studied CO in cc-pVTZ basisDunning (1989) at bond-distances between and Å. There, we used amfX2C (XAMFI) correctionsKnecht et al. (2022) for scalar-relativistic and spin-orbit two-electron picture-change effects arising within the X2C Hamiltonian framework.Knecht et al. (2022) HF orbitals for a given point on the curve served as an initial guess for the SCF calculation at the next point with increasing distances.
Then we calculated the MP2 frozen natural orbitalsYuan, Visscher, and Gomes (2022) (with a cutoff on occupation number) on top of which the final CC calculation was performed.
We performed two correctness checks with a nonrelativistic Hamiltonian. For the first test, we chose H2O in a 3-21G basisBinkley, Pople, and Hehre (1980) at the equilibrium internuclear distance of Å and H-O-H angle of . The second test is with LiH in 6-31G basisDill and Pople (1975); Ditchfield, Hehre, and Pople (1971) at the internuclear distance of Å. Both the CC correctness tests were performed with respect to the HF reference.
All the CC calculations were performed in uncontracted basis sets. When referring to sizes of occupied and virtual spaces, we use the notation of CAS(), where is the number of electrons and is the number of Kramers pairs active in a coupled cluster calculation (occupied+virtual). One Kramers pair corresponds to two spinors related by time reversal symmetry. In all calculations, the choice of occupied and virtual spaces respects the rising orbital energies or the occupation numbers in case of MP2. See the used software in section IV.1.
@@ IV.1 Correctness: H2O and LiH
To demonstrate the correctness, we compare ground state coupled cluster energies from tenpi (present work) with the established MRCC packageKállay et al. (2020); mrc (2024); Kállay and Surján (2001); Kállay, Gauss, and Szalay (2003) and with existing code in DIRACVisscher et al. (2024); Saue et al. (2020) where possible, i.e. either the initial DIRAC RelCC moduleVisscher, Lee, and Dyall (1996); Shee, Visscher, and Saue (2016) or the recent codegen (mb-autogen).Gomes and Lyakh (2023) The comparison was performed on two small systems H2O (Table 3) and LiH (Table 4) and shows an agreement within the convergence thresholds.
| method | code | convergence555the final error shown in the run | Total energy [EH] |
@@
| CCSD | DIRAC tenpi444using the single node TAL-SH implementation | 0.5E-09333norm of the residual vector | -75.67 2947 6512 |
@@
| method | code | convergence444the final error shown in the run | Total energy [EH] |
@@
IV.2 Benchmark: UF6 CCSD strong scaling
To benchmark the parallel scaling, we compare the performance of CCSD generated by tenpi with hand-tuned code in the ExaCorr module of DIRAC and with generated code from the codegen (mb-autogen) generator.Gomes and Lyakh (2023) We perform a strong-scaling benchmark on UF6. In all cases, the correct total energy of is retrieved with residual norm threshold less than . The parallel speedup for strong scaling benchmark is calculated in a standard way as a walltime ratio for processing units (in this case we consider entire nodes). If is not available due to large memory requirements, the timing per node with the miniminal number of nodes is used to estimate it.
The benchmark was performed on the Summit supercomputer of the Oak Ridge Leadership Computing Facility using the GPU implementation of the ExaTENSOR library. Each Summit node is equiped with 6 GPUs (see Table 2). We used up to 300 GPUs simultaneously.
As shown in Figure 5, the code scales well until about 20 nodes for UF6. There, the curve leaves its approximately linear behaviour. This is expected as the internode communication overhead increases with the number of nodes. This can be modelled via Amdahl’s lawAmdahl (2013) for strong scaling with the timing proportion of ‘serial code’ at when the processes wait for communication during tensor operations and for synchronization barriers which are enforced after each tensor operation.
As expected, the automatically generated code is slightly less efficient than the hand-tuned code. The timings in Figure 6 show a fixed-factor slowdown of about 30% for tenpi, whereas about 80% for codegen. The former is a good result when compared with ORCA, where 100% timing overhead was a practical rule of thumb threshold for acceptance of generated code. Lechner et al. (2024)

### v2

@@ II THEORY AND IMPLEMENTATION
II.1 Coupled cluster amplitude equations and their generation
@@
where is an excitation operator, and the corresponding cluster amplitude tensor. in our case denotes the Hartree–Fock (HF) reference.
@@
We use and to denote occupied and virtual orbitals respectively.
After defining the similarity transformed normal-ordered Hamiltonian as , we can write down the CC energy equations
@@
The explicit equation terms can then be derived using different techniques. In our implementation, we chose the diagrammatic scheme of Kállay and SurjánKállay and Surján (2001) (see the reasoning in Section II.2). This general-order coupled cluster derivation scheme is based on a string representation of CC diagrams.
As depicted in Fig. 2, an integer string of length 13 is used to represent a diagram. Two diagrams are equivalent up to a sign if their strings are the same. There is a simple algorithm depicted in Fig. 3 that uses this representation to generate distinct diagrams for each excitation level of the CC Eqs. (3).
To find the explicit equation terms, tenpiBrandejs and Saue (2024) translates the integer strings into line graph representation, i.e. an edge list of particle and hole lines connecting operators and external bra indices of Eqs. (3). Such a representation makes the application of coupled cluster interpretation rules similar to when done by hand. All used high-level representations are summarized in Fig. 5 and explained in Sections II.2 and II.3.
II.2 Design decisions
The initial goal of the tenpi project was to create a programming environment for relativistic coupled cluster which separates science from the computational platform by getting tensor developments under the hood. This has three aspects:
@@ II.2 Design decisions
Automatic treatment of tensor symmetries, such as index-permutation symmetry and block-sparsity (for systems with spatial symmetry).
@@
The design choices are limited by practical considerations. The toolchain should not be restricted to a given flavor of the coupled cluster method, but should be written in a general way and be able to go to high orders of the theory. It should allow manipulations of the method by PhD and Master students of quantum chemistry. This led to the decision to base the formula generator on Kucharski-Bartlett diagrams,Shavitt and Bartlett (2009); Kucharski and Bartlett (1992) since diagrams are visual and thus more accessible for students than the other CC derivation schemes listed in section I.1. The ease of use for target users is one of the reasons why we chose python to implement the generator, aside from the fact that it is suitable for text processing and symbolic operations. The string representation of CC diagrams of Kállay and SurjánKállay and Surján (2001) was chosen because it clearly distinguishes equivalent diagrams and can be used to generate distinct connected diagrams (see section II.1) in an intuitive way.Kállay and Surján (2001)
@@
The diagram strings are next translated into directed graphs, by storing a list of edges (adjacency matrix). This representation makes the implementation of the CC interpretation rules readable. Advanced users can define a custom set of rules (customizability) or for instance extend the Diagram class to support multiple interaction operators and other generalizations, without having to change the rest of the code (modularity).
@@
To follow closely the hand-typeset style of the pedagogical texts of Refs. Crawford and Schaefer, 2000 and Shavitt and Bartlett, 2009, tenpi prints the diagrams in PDF format using the CCDiag LaTeXpackage.Kats (2024) During the development of higher-order CC methods this visual representation proved itself invaluable for debugging. Overall, the emphasis on intuitive, readable, modular and visual implementation makes problems directly visible and modules unit-testable (see section II.3).
@@
The workflow of the tenpi code generator is depicted in Fig. 4 and corresponding high-level representations in Fig. 5. A key step is to optimize the expressions obtained. Since writing a global optimizer from scratch is beyond the scope of this work, we interfaced our code with OpMin, which was chosen over GristmillSong, Henderson, and Scuseria (2022) due to a favorable performance for higher-order CC.Lai et al. (2012) However, even though the heuristics of OpMin is reported to stay within of the global minimum FLOP-count, the number of intermediates it generates tends to be too costly in terms of memory. Higher-order CC tensors in our calculations cannot be stored on disk due to prohibitive I/O costs. Our approach is to distribute them among the CPU and GPU RAM memory of the nodes. This comes with a limitation in the amount of available distributed memory due to communication overhead growing quadratically with the number of nodes in the round-robin distribution scheme used (see below). Therefore, we implemented our own secondary optimization, which rolls back part of the intermediates by plugging them back in their respective terms. The number of intermediates is reduced by about a factor of three at the expense of increasing the FLOP count (see Fig. 6). This approach is justified as long as the calculation stays communication-dominated. We do not reach a perfect balance, which would require setting up a communication-based cost-function. Optimization of this aspect is the subject of ongoing work.
@@
During the optimization, we originally represented contractions using the SymPy packageMeurer et al. (2017) but later we turned to the use of a custom representation of tensor contractions for greater flexibility (Contraction class, see section II.3). Even though SymPy is widely used, its current user interface is cumbersome for higher-order tensors.
@@
As we aim for scalability on modern GPU exascale machines, we turned to the use of a tensor library to execute the final tensor kernels (ExaTENSOR, TAL-SH). ExaTENSOR was chosen from the limited list of available distributed GPU tensor libraries (see section I.3) due to its excellent performance for relativistic CCSD.Pototschnig et al. (2021) In the present work, the performance is reproduced and demonstrated to scale to up to 1200 GPUs by automatically generated code (see Fig. 9).
II.3 Implementation
An equation term is implemented using the ‘Contraction’ class, which includes a list of tensors, a scalar factor and optionally permutation operators and an output tensor. Contraction class instances are collected in a list, which forms the left-hand side of Eqs. (3).
Tensor class includes a list of upper and lower indices. Each index is an Index class instance. Index class has its type (e.g. occupied, virtual, active virtual), a convention according to which it is printed (e.g. ; ) and then a number corresponding to the order in which the specific character comes in the class (e.g. 1, 2, 1). Types and conventions are easily customizable by modifying a single location in the code, which greatly facilitates the implementation of methods and symmetries with multiple index spaces or tiling.Mutlu et al. (2023) Index, Tensor and Contraction class are fully compatible with smart functionality to account for block-sparsity in tensors and to keep track of index permutation symmetry in intermediates explicitly. Work is in progress to include this functionality in the production code. A set of equations can switch between conventions with a single command. This is important because tenpi supports multiple input and output formats to interface with ExaTENSOR, TAL-SH, OpMin, NumPy, including a possibility to input equations by hand (see Fig. 5). For the sake of modularity, all these codes are integrated by extending one of two predefined interfaces: GetCodeInterface or ParseCodeInterface. These specify how to implement a new format of output or input respectively. Thanks to this design it is straightforward to interface tenpi with a new program.
@@ II.3 Implementation
| OpMin | 0.03 s | 40 s | 22 min | 7 h 10 min | 4 days 1 h |
| 2nd optimization | 0.02 s | 0.1 s | 16 s | 1 h 9 min | 17 days |
@@
The derivation step of tenpi is very fast, see timings in Tab. 1. We have verified by comparing with Ref. Kállay and Surján, 2001 that tenpi generates the correct number of diagrams for each of the CC Eqs. (3) up to excitation level 20. Even though it generates a highly-parallel code, tenpi itself is written as sequential python. Despite this, the most demanding step of global intermediate optimization still performs quite well.
Aside from amplitude equations, tenpi can already generate matrix elements (see Figs. 3, 5, cf. SMITHShiozaki et al. (2008); Shiozaki (2017)) and is currently being extended to generate response density matrices.
When going from CCD and CCSD levels during the development, an error in energy appeared. It was found quickly thanks to the visual representation of diagrams: For the CC diagram interpretation rules to yield a correct sign, e.g. for the case of amplitude-equation projected onto , the line starting in has to end in , and not in . This rule is often assumed implicitly.Shavitt and Bartlett (2009); Crawford and Schaefer (2000) Transition from CCSDT and CCSDTQ proved to be more peculiar, due to a mistake in an algorithm for unrolling index permutations. The problem was unraveled once a brute force unit test was written for the corresponding module.
@@ III.1 Machines used
The calculations were performed on the machines listed in Table 2 which includes their technical specification. When ExaTENSOR is applied for coupled cluster,Pototschnig et al. (2021) the most relevant parameters for the performance are the interconnect bandwidth, RAM (random-access) and HBM (high-bandwidth) memory, as well as GPU specifications. Over the two supercomputer generations represented in Table 2 the bandwidth and the HBM have grown about fourfold, while RAM has basically stagnated. Despite the rapid development of GPU processing power, the table shows that the evolution of other important parameters has been relatively slow over the last 5 years. Recently, large memory CPU nodes have emerged which go against this trend, e.g. Intel Xeon 6900P with MRDIMM technology (Multiplexed Rank Dual In-line Memory Module) and 3 TB of fast memory as announced for an HPC platform in Japan.NEC (2024)
We would like to bring to the attention of the reader that one should not expect an easy installation and tuning of distributed memory tensor libraries, as it mostly requires GPU and MPI experts. ExaTENSOR features cutting-edge parallelization strategy with multithreading, nested OpenMP threads, GPU-aware MPI and one-sided communication, which proved itself difficult to support for some HPC platforms. For instance on Frontier, an internal resource starvation within Cray MPICHHewlett Packard Enterprise (2023) (HPE) causes certain MPI requests to be blocked, and the calculation hangs as a result. The engineers from the supplier company (HPE) are currently working on the issue.
@@ III.2 Relativistic Hamiltonian
The state-of-the-art post-HF four-component relativistic calculations employ the no-pair approximation.Sucher (1980); Almoukhalalati et al. (2016) This approximation is well-substantiated for applications of chemical scale. Formally, it yields a Hamiltonian analogous to the non-relativistic case
where the indices only run over the positive-energy spinors that span the one-electron basis.
In the present work we employ, unless otherwise stated, the exact two-component (X2C) Hamiltonian,Jensen (2005); Kutzelnigg and Liu (2005); Iliaš and Saue (2007); Liu and Peng (2009) where the summations in Eq. (4) are limited to positive energy orbitals by construction. Relativistic two-electron picture-change corrections were added either using the AMFI packageHess et al. (1996); Schimmelpfennig or using the recently implemented amfX2C (XAMFI) correction.Knecht et al. (2022) The Gaunt/Breit terms are not included in the present work.
@@
For the benchmarks, we used the UF6 molecule in dyall.v2z basisDyall (2023) with the X2C Hamiltonian, and bond distance of Å. In the corresponding ground-state energy CC calculations, we used a HF reference. Despite the high Oh symmetry, there is no related performance gain in the CC calculations as the implementation does not support symmetry yet.
Furthermore, we studied CO in cc-pVTZ basisDunning (1989) at bond-distances between and Å. There, we used the amfX2C correction for two-electron scalar-relativistic and spin-orbit picture-change effects arising within the X2C Hamiltonian framework.Knecht et al. (2022)
For the cases where the full virtual space is not used, we then calculated the MP2 frozen natural orbitalsYuan, Visscher, and Gomes (2022) (with a cutoff on occupation number) on top of which the final CC calculation was done.
We performed two correctness checks with a nonrelativistic Hamiltonian. For the first test, we chose H2O in a 3-21G basisBinkley, Pople, and Hehre (1980) at the equilibrium internuclear distance of Å and H-O-H angle of . The second test is with LiH in 6-31G basisDill and Pople (1975); Ditchfield, Hehre, and Pople (1971) at the internuclear distance of Å. Both CC correctness tests were performed with respect to the HF reference.
All the CC calculations were performed in uncontracted basis sets. When referring to sizes of occupied and virtual spaces, we use the notation of AS(), where is the number of electrons and is the number of Kramers pairs active in a coupled cluster calculation (occupied+virtual). One Kramers pair corresponds to two spinors related by time reversal symmetry. In all calculations, the choice of occupied and virtual spaces respects the rising orbital energies or the occupation numbers in the case of MP2. See the used software in Section IV.1.
@@ IV.1 Correctness: H2O and LiH
To demonstrate the correctness, we compare ground state coupled cluster energies from tenpi (present work) with the established MRCC packageKállay et al. (2020); mrc (2024); Kállay and Surján (2001); Kállay, Gauss, and Szalay (2003) and with existing handwritten code in DIRACVisscher et al. (2024); Saue et al. (2020) where possible, i.e. either with the initial DIRAC RelCC moduleVisscher, Lee, and Dyall (1996) or with the more recent massively parallel ExaCorr module.Pototschnig et al. (2021) The correctness check was performed on two small systems: H2O (Table 3) and LiH (Table 4) and shows an agreement within the convergence error.
| method | code | convergence555the final error shown in the run | Total energy [] |
@@
| CCSD | DIRAC tenpi444using the TAL-SH library | 0.5E-09333norm of the residual vector | -75.67 2947 6512 |
@@
| method | code | convergence444the final error shown in the run | Total energy [] |
@@
IV.2 Benchmark: UF6 CCSD
To benchmark the parallel scaling, we compare the performance of CCSD generated by tenpi with hand-tuned code in the ExaCorr module of DIRAC and with generated code from the codegen (mb-autogen) generator.Gomes and Lyakh (2023) First we perform a strong-scaling benchmark on UF6 with fixed problem size. In all cases, the correct total energy of is retrieved with residual norm less than . The parallel speedup for strong scaling benchmark is calculated in a standard way as a walltime ratio for processing units (in this case we consider entire nodes). If is not available due to large memory requirements, the timing per node with the minimal number of nodes is used to estimate it.
The benchmark was performed on the Summit supercomputer of the Oak Ridge Leadership Computing Facility (OLCF) using the GPU implementation of the ExaTENSOR library. Each Summit node is equipped with 6 GPUs (see Table 2). We used up to 300 GPUs simultaneously for strong scaling.
As shown in Fig. 7, the code scales well until about 20 nodes for UF6. There, the curve leaves its approximately linear behavior. This is expected for fixed problem size. As the number of workers increase, they start to suffer from work starvation, which eventually causes the overhead from the unnecessarily increased internode communication to dominate. We chose to model this using the Amdahl’s lawAmdahl (2013) for strong scaling
where is the number of nodes, and are the proportions of serial and parallel code. From fitting, we found , corresponding to when the processes wait for communication during tensor operations and for synchronization barriers which are enforced after each tensor operation.
As expected, the automatically generated code is slightly less efficient than the hand-tuned code. The timings in Fig. 8 show a fixed-factor slowdown of about 30% for tenpi, whereas about 80% for codegen. The former is a good result when compared with ORCA, where 100% timing overhead was a practical rule of thumb threshold for acceptance of generated code. Lechner et al. (2024)
To investigate the influence of the internode communication, we then moved to analyze the weak scaling of the code. Using the same system and method as previously, we varied the size of virtual space such that the number of operations per node is kept fixed, based on expected CCSD scaling of with the number of occupied and of virtual spinors. The weak scaling efficiency is again calculated as . The ideal weak scaling efficiency would be a constant 1.
As demonstrated in Fig. 9, the code exhibits a relatively constant weak scaling behavior to up to 200 nodes, meaning that it can be used efficiently on up to 1200 GPUs. This is a pleasing result for a code without hand-tuning. We attribute most fluctuations in the graph to differing appropriateness of fixed block size of elements used in case of CCSD when distributing the tensors, where is the number of indices. Some virtual sizes come with smaller number of padding zeroes than others. Moreover, the handwritten code has different intermediates than tenpi and the advantage can differ for different space sizes. One could profile and hand-tune the generated code based on several test runs, but such non-systematic intervention would go against the design principles. Instead, using measured walltimes to improve the cost model in the optimization step is a subject of further study.